# Simulating Optical Coherent Nonlinear Response for High-Intensity Excitation

### v1

@@ III Nonlinear Signal saturation
We use a collinear 2D Coherent Spectroscopy (2DCS) experiment, which is similar to the commonly-used acousto-optic modulator (AOM) based experiments Martin and Cundiff (2018); Nardin et al. (2013); Tekavec, Lott, and Marcus (2007), and is briefly described here. The excitation pulses are derived from the output of an optical parametric amplifier (OPA) system with a repetition rate of 50 kHz. The output of the OPA is centered at 755 nm and has a full-width at half-maximum (FWHM) bandwidth of $\sim 20$ nm, which results in a transform-limited pulse duration of $\sim 60$ fs. These pulses are then split into four different paths (A, B, C and D), and are tagged with unique radio frequencies – $\Omega_{A}$=78.858 MHz, $\Omega_{B}$=78.595 MHz, $\Omega_{C}$=79.597 MHz, and $\Omega_{D}$ =79.304 MHz through phase modulation in the AOMs. The drive frequencies for the AOMs are generated through a direct digital synthesizer and are delivered to the AOMs using high-voltage drivers. These pulses are recombined to obtain the excitation pulses with a specific time-ordering that is shown in Fig. 2. The delays between the consecutive pulses – $\tau_{1}$, $\tau_{2}$, and $\tau_{3}$ can be scanned using motorized delay stages. The FWHM bandwidth of excitation pulses is reduced to $\sim 7$ nm using a grating-based pulse shaper. The narrow spectrum of the excitation pulse, shown in Fig. 3(a), ensures that only the QW excitons are excited. The excitation pulses with collinear polarization are focused onto the sample using a lens with a focal length of 20 cm, resulting in a spot size with a $1/e^{2}$ radius of $\sim 100\>\mu$m. All the excitation pulses have equal intensity, which is controlled using a combination of a half-wave plate and a polarizer. A continuous wave (CW) laser with a wavelength of 1064 nm is used for passive stabilization of the pulse delays. The CW laser traverses the same optical path as the excitation laser and is, thus, modulated at the same AOM frequencies. Two photodiodes measure the beat notes arising from the interference of the CW laser from different paths – $\Omega_{AB}=263$ kHz for the A-B path and $\Omega_{CD}=293$ kHz from the C-D path. These beat notes provide the reference $\Omega_{FWM}=\Omega_{CD}-\Omega_{AB}=30$ kHz to measure the rephasing FWM signal using a lock-in amplifier. The FWM signal is measured as a function of variable delays $\tau_{1}$ and $\tau_{3}$ while keeping delay $\tau_{2}$ fixed at 0.7 ps. Delay $\tau_{1}$ is scanned in discrete steps, while delay $\tau_{3}$ is scanned continuously to reduce the data-aquisition time. The generation of the phase-stabilized reference, lock-in detection, and scanning of the time delays is performed using a customized Bigfoot electronics from MONSTR Sense that uses a field-programmable gate array (FPGA) Huang et al. (2023); Połczyńska et al. (2023). The electronics system generates a virtual reference signal corresponding to a wavelength close to the exciton resonance, which reduces the phase fluctuations on the lock-in detected signal significantly.
@@
We measured rephasing spectra as a function of intensity of excitation pulses. The intensity-dependent peak amplitude, which is calculated using the procedure outlined in Appendix B, is shown as a log-log plot in Fig. 4. Figure 4 also shows the expected $\chi^{(3)}$-dependence in dashed line. At low intensities (shaded region in Fig. 4), the signal amplitude follows this behavior, which is roughly in agreement with prediction of third-order perturbation theory. However, as the average power per excitation pulse exceeds $\sim 0.7\mu$W, the data deviates from the expected $\chi^{(3)}$-dependence and saturates. This behavior suggests that at high intensities, third-order perturbation theory does not accurately describe the FWM signal amplitude, and higher-order nonlinear signals contribute significantly to the total signal.
We reproduce the experimental results using simulations based on the phase-cycling method discussed in Sec. II. We model the excitonic system as a two-level system, as shown in Fig. 1(a). We consider the sum of the electric fields of four excitation pulses, as shown in Fig. 2, and numerically solve the OBEs in Eqs. 3 as a function of delays $\tau_{1}$ and $\tau_{3}$. The most appropriate and economical $3\times 3\times 3\times 1$ phase-cycling scheme was used to acquire the rephasing signal as a function of the variable delays. The 2D rephasing spectrum was then obtained by performing a Fourier transform with respect to both the time delays. Various rephasing spectra were simulated with increasing pulse areas of each excitation pulse. The log-log plot of the maximum amplitude of each spectrum against the pulse area, indicated by the solid blue line in Fig. 4, shows an excellent match with the data; we use an arbitrary scaling factor to relate the pulse area used in the simulations to the square-root of the power used in the experiments.
Figure 4: Log-log plot of the amplitude of the experimental rephasing spectrum (green circles) as a function of the power of excitation pulses, and the amplitude of the simulated rephasing spectrum (solid blue line) versus the pulse area of the pulses. The data point encased in a box corresponds to the 2D spectra shown in Fig. 3(b). The dashed red line indicates the expected $\chi^{(3)}$ dependence. The shaded region marks the $\chi^{(3)}$ regime. The lower X-axis represents the pulse area for each point of the simulated 2D spectra, while the upper X-axis corresponds to the square-root of the power of the excitation pulses in the experiments. The simulation data points are normalized by the peak amplitude at maximum pulse area and the experimental data points were rescaled along both axes to match the scale of the simulation data.
@@ IV.1 Higher-order features
Figure 5: An energy-level diagram for the exciton-biexciton diamond-type system depicting the ground state |$\ket{G}$, the lowest-energy orthogonal exciton states |$\ket{H}$ and |$\ket{V}$, and the bound biexciton state |$\ket{B}$, which is red-shifted from the unbound two-exciton state by the biexciton binding energy $\Delta_{XX}$.
We begin with a study Moody et al. (2013) where $\chi^{(5)}$ signal from the biexciton state in self-assembled InAs QDs was spectrally isolated in a 2DCS experiment. This study used cross-linearly polarized (HVVH) excitation sequences to selectively excite the biexcitonic state in neutral QDs. The FWM signal was measured in the phase-matched direction $\mathbf{k}_{S}=-\mathbf{k}_{A}+\mathbf{k}_{B}+\mathbf{k}_{C}$. At low excitation intensities, i.e. withing the $\chi^{(3)}$ regime, the 2D spectrum displayed a single biexciton peak. However, at higher intensities, an additional biexciton peak emerged, which was attributed to the $\chi^{(5)}$ nonlinear response. Double-sided Feynman diagrams, with contributions up to fifth-order excitation were used to explain the origin of this peak.
@@
We model the exciton-biexciton system as a diamond-type system shown in Fig. 5, and the OBEs were modeified accordingly Suzuki et al. (2016). We incorporated polarization of excitation pulse in the electric fields used in these equations to ensure a cross-linear excitation scheme. For simplicity, the trion states, inhomogeneity, and the fine-structure splitting were ignored since they are not relevant for the results discussed here. Biexciton binding energy $\Delta_{XX}=3.3$ meV was used. Similar to the previous section, the $3\times 3\times 3\times 1$ phase cycling method was chosen to simulate the rephasing signal satisfying the phase matching condition $\phi_{sig}=-\phi_{A}+\phi_{B}+\phi_{C}-\phi_{D}$. We varied the pulse area of the excitation pulses in our simulation. At low intensities, consistent with the $\chi^{(3)}$ regime, the simulation results showed a single biexciton peak XXL, as depicted in Fig. 6(a). As the pulse intensity increased, the simulated spectrum in Fig. 6(b) revealed the emergence of an additional biexciton peak XXU. These peaks are shifted from the diagonal along the excitation and emission axes by the biexciton binding energy, which is consistent with the experiment. As was done in the original study, a comparison of the excitation-intensity dependence of the peak amplitudes highlights the distinct perturbative order contributions of $\chi^{(3)}$ and $\chi^{(5)}$ to these peaks, as shown in Fig. 7.
@@ IV.2 Switching between nonlinear signals
We next focus on a study Fras et al. (2016) that reported conversion of FWM signal to SWM signal in a single self-assembled InAs QD, showcasing a novel coherent-control mechanism in single quantum emitters. The researchers utilized three excitation pulses with wavevectors $\mathbf{k_{i}}$, where $i\in{1,2,3}$, to resonantly drive a single InAs quantum dot. The first two pulses create a FWM transient, which evolves freely until the arrival of the third pulse. This third pulse converts the FWM signal $\xi_{F}$ emitted in the direction $\mathbf{k}_{F}=2\mathbf{k}_{2}-\mathbf{k}_{1}$ into SWM signal $\xi_{S}$ emitted in direction $\mathbf{k}_{S}=\mathbf{k}_{1}-2\mathbf{k}_{2}+2\mathbf{k}_{3}$, achieving up to unity conversion efficiency. This conversion process depends on the delay and pulse area of the third pulse, aligning with theoretical model derived by assuming excitation with $\delta-$function pulses:
@@
The optimal phase cycling methods for the signals corresponding to $\phi_{F}$ and $\phi_{S}$ were $3\times 6\times 3\times 1$ and $3\times 5\times 4\times 1$, respectively. However, we used the $3\times 6\times 4\times 1$ phase cycling method, which allowed us to isolate both signals simultaneously using the same phase combination of excitation pulses. This phase-cycling scheme is the optimal choice for our purpose. As an added advantage, using the same phase-cycling scheme for both the FWM and SWM signals allows a direct comparison of their amplitudes. This approach is also relevant for experimental settings where the same phase combination can be used to simultaneously measure both signals. We model a single QD as a 2LS and used the OBEs incorporating all decay terms similar to those in the discussed study. Our simulation results, depicted in Fig. 8 for $\theta_{3}=\pi$, show similar observations to those described in the study. During the time interval between the arrival of the second and third excitation pulses, the FWM signal decays at the population decay rate (0.01 meV) while the SWM signal amplitude remains zero. Consistent with the analytical calculations, we observed that as the third pulse arrived ($\tau_{3}>0$ ps), the total FWM signal is converted to the SWM signal, and the FWM signal value dropped to zero.
@@ IV.3 Photon echo from high-intensity excitation
As a final example, we look at a study that investigated the effect of high-intensity excitation pulses on the photon-echo transients in an ensemble of (In,Ga)As/GaAs QDs Poltavtsev et al. (2016). In a two-pulse FWM experiment, the photon-echo signal peaks after the second excitation pulse at a delay equal to the delay between the excitation pulses. In Ref. Poltavtsev et al., 2016, experiments performed with high-intensity excitation pulses show that the photon-echo transients do not necessarily follow this behavior. Solutions of the Lindblad equation assuming rectangular temporal profile of the excitation pulses were used to model the experimental findings and infer that significant modifications in the photon-echo transients occur when the spectral width of the inhomogeneous ensemble is comparable to those of the excitation pulses.
@@
The above results explicitly demonstrate that the our calculations also work for systems with inhomogeneous broadening. As in Sec. IV.2, we did not make any simplifying assumptions about the pulse shape, as was done in the original work Poltavtsev et al. (2016). Collectively, the results presented in this section highlight the effectiveness of the phase-cycling method to calculate the exact nonlinear response even for high-intensity excitation pulses.
@@ V Conclusion
In this work we have presented a comprehensive study on the application of the phase-cycling method to simulate coherent nonlinear signal for high-intensity excitation pulses. We started with a generalized description of the phase-cycling method that can be applied to an arbitrary NWM signal. Following this, we used the phase-cycling method to reproduce the saturation of the FWM signal as the excitation pulse intensity is increased. Finally, we reproduced three previously-reported experimental results on measuring coherent nonlinear response of QDs using high-intensity excitation pulses.
@@
Appendix A Optimized phase-cycling scheme

### v2

@@ III Nonlinear Signal saturation
We use a collinear 2DCS experiment, which is similar to the commonly-used acousto-optic modulator (AOM) based experiments Martin and Cundiff (2018); Nardin et al. (2013); Tekavec, Lott, and Marcus (2007), and is briefly described here. The excitation pulses are derived from the output of an optical parametric amplifier (OPA) system with a repetition rate of 50 kHz. The output of the OPA is centered at 755 nm and has a full-width at half-maximum (FWHM) bandwidth of $\sim 20$ nm, which results in a transform-limited pulse duration of $\sim 40$ fs. These pulses are then split into four different paths (A, B, C and D), and are tagged with unique radio frequencies – $\Omega_{A}=78.858$ MHz, $\Omega_{B}=78.595$ MHz, $\Omega_{C}=79.597$ MHz, and $\Omega_{D}=79.304$ MHz through phase modulation in the AOMs. The drive frequencies for the AOMs are generated through a direct digital synthesizer and are delivered to the AOMs using high-voltage drivers. These pulses are recombined to obtain the excitation pulses with a specific time-ordering that is shown in Fig. 2. The delays between the consecutive pulses – $\tau_{1}$, $\tau_{2}$, and $\tau_{3}$ can be scanned using motorized delay stages. The FWHM bandwidth of excitation pulses is reduced to $\sim 6$ nm using a grating-based pulse shaper. The narrow spectrum of the excitation pulse, shown in Fig. 3(a), ensures that only the QW excitons are excited. The excitation pulses with collinear polarization are focused onto the sample using a lens with a focal length of 20 cm, resulting in a spot size with a $1/e^{2}$ radius of $\sim 100\>\mu$m. All the excitation pulses have equal intensity, which is controlled using a combination of a half-wave plate and a polarizer. A continuous wave (CW) laser with a wavelength of 1064 nm is used for passive stabilization of the pulse delays. The CW laser traverses the same optical path as the excitation laser and is, thus, modulated at the same AOM frequencies. Two photodiodes measure the beat notes arising from the interference of the CW laser from different paths – $\Omega_{AB}=263$ kHz from the A-B path and $\Omega_{CD}=293$ kHz from the C-D path. These beat notes provide the reference $\Omega_{FWM}=\Omega_{CD}-\Omega_{AB}=30$ kHz to measure the rephasing FWM signal using a lock-in amplifier. The FWM signal is measured as a function of variable delays $\tau_{1}$ and $\tau_{3}$ while keeping delay $\tau_{2}$ fixed at 0.7 ps. Delay $\tau_{1}$ is scanned in discrete steps, while delay $\tau_{3}$ is scanned continuously to reduce the data-aquisition time. The generation of the phase-stabilized reference, lock-in detection, and scanning of the time delays is performed using a customized Bigfoot electronics from MONSTR Sense that uses a field-programmable gate array (FPGA) Huang et al. (2023); Połczyńska et al. (2023). The electronics system generates a virtual reference signal corresponding to a wavelength close to the exciton resonance, which reduces the phase fluctuations on the lock-in detected signal significantly.
@@
We measured rephasing spectra as a function of intensity of excitation pulses. The intensity-dependent peak amplitude, which is calculated using the procedure outlined in Appendix C, is shown as a log-log plot in Fig. 4. Figure 4 also shows the expected $\chi^{(3)}$-dependence in dashed line. At low intensities (shaded region in Fig. 4), the signal amplitude follows this behavior, which is roughly in agreement with prediction of third-order perturbation theory. However, as the average power per excitation pulse exceeds $\sim 0.7$ $\mu$W, the data deviates from the expected $\chi^{(3)}$-dependence and saturates. This behavior suggests that at high intensities, third-order perturbation theory does not accurately describe the FWM signal amplitude, and higher-order nonlinear signals contribute significantly to the total signal.
Figure 4: Log-log plot of the amplitude of the experimental rephasing spectrum (green circles) as a function of the power of excitation pulses, and the amplitude of the simulated rephasing spectrum (solid blue line) versus the pulse area of the excitation pulses. The data point encased in a box corresponds to the 2D spectra shown in Fig. 3(b). The dashed red line indicates the expected $\chi^{(3)}$ dependence. The shaded region marks the $\chi^{(3)}$ regime. The lower X-axis represents the pulse area for each point of the simulated 2D spectra, while the upper X-axis corresponds to the square-root of the power of the excitation pulses in the experiments. The simulation data points are normalized by the peak amplitude at maximum pulse area and the experimental values were rescaled to match the normalized simulation data.
We reproduce the experimental results using simulations based on the phase-cycling method discussed in Sec. II. We model the excitonic system as a two-level system, as shown in Fig. 1(a). We consider the sum of the electric fields of four excitation pulses, as shown in Fig. 2, and numerically solve the OBEs in Eqs. 3 as a function of delays $\tau_{1}$ and $\tau_{3}$. The most appropriate and economical $3\times 3\times 3\times 1$ phase-cycling scheme was used to acquire the rephasing signal as a function of the variable delays. The 2D rephasing spectrum was then obtained by performing a Fourier transform with respect to both the time delays. Various rephasing spectra were simulated with increasing pulse areas of each excitation pulse. The log-log plot of the maximum amplitude of each spectrum against the pulse area, indicated by the solid blue line in Fig. 4, shows an excellent match with the data for $\mu\approx 630$ D, which is consistent with previously reported value for QW excitons Christmann et al. (2011).
@@ IV Phase-Cycling Simulations of Previous Experimental Works
Figure 5: An energy-level diagram for the exciton-biexciton diamond-type system depicting the ground state |$\ket{G}$, the lowest-energy exciton states |$\ket{H}$ and |$\ket{V}$, and the bound biexciton state |$\ket{B}$, which is red-shifted from the unbound two-exciton state by the biexciton binding energy $\Delta_{XX}$. The |$\ket{H}$ and |$\ket{V}$ exciton states are excited by light with orthogonal linear polarizations H and V, which are indicated by the solid and dashed arrows, respectively.
@@ IV.1 Higher-order features
We begin with a study Moody et al. (2013) where $\chi^{(5)}$ signal from the biexciton state in self-assembled InAs QDs was spectrally isolated in a noncollinear 2DCS experiment. This study used cross-linearly polarized (HVVH) excitation sequences to selectively excite the biexcitonic state in neutral QDs. The FWM signal was measured in the phase-matched direction $\mathbf{k}_{S}=-\mathbf{k}_{A}+\mathbf{k}_{B}+\mathbf{k}_{C}$. At low excitation intensities, i.e. within the $\chi^{(3)}$ regime, the 2D spectrum displayed a single biexciton peak. However, at higher intensities, an additional biexciton peak emerged, which was attributed to the $\chi^{(5)}$ nonlinear response. DSFDs, with contributions up to fifth-order excitation were used to explain the origin of this peak.
@@
We model the exciton-biexciton system as a diamond-type system shown in Fig. 5, and the OBEs were modified accordingly Suzuki et al. (2016)(see Appendix D for the detailed equations). We incorporated polarization of excitation pulse in the electric fields used in these equations to ensure a cross-linear excitation scheme. For simplicity, the trion states, inhomogeneity, and the fine-structure splitting were ignored since they are not relevant for the results discussed here. Biexciton binding energy $\Delta_{XX}=3.3$ meV was used. Similar to the previous section, the $3\times 3\times 3\times 1$ phase cycling method was chosen to simulate the rephasing signal satisfying the phase matching condition $\phi_{sig}=-\phi_{A}+\phi_{B}+\phi_{C}-\phi_{D}$. We varied the pulse area of the excitation pulses in our simulation. At low intensities, consistent with the $\chi^{(3)}$ regime, the simulation results showed a single biexciton peak XXL, as depicted in Fig. 6(a). As the pulse intensity increased, the simulated spectrum in Fig. 6(b) revealed the emergence of an additional biexciton peak XXU. These peaks are shifted from the diagonal along the excitation and emission axes by the biexciton binding energy, which is consistent with the experiment. As was done in the original study, a comparison of the excitation-intensity dependence of the peak amplitudes highlights the distinct perturbative order contributions of $\chi^{(3)}$ and $\chi^{(5)}$ to these peaks, as shown in Fig. 7.
@@ IV.2 Switching between nonlinear signals
We next focus on a study Fras et al. (2016) that reported conversion of FWM signal to SWM signal in a single self-assembled InAs QD, showcasing a novel coherent-control mechanism in single quantum emitters. The researchers employed three collinear excitation pulses, each uniquely tagged with a frequency $\Omega_{i}$ ($i\in{1,2,3}$) using AOMs, to resonantly drive a single InAs quantum dot. The first two pulses create a FWM transient, which evolves freely until the arrival of the third pulse. This third pulse converts the FWM signal $\xi_{F}$, detected at the heterodyne frequency $\Omega_{F}=2\Omega_{2}-\Omega_{1}$ into SWM signal $\xi_{S}$, detected at the heterodyne frequency $\Omega_{S}=\Omega_{1}-2\Omega_{2}+2\Omega_{3}$, achieving up to unity conversion efficiency. This conversion process depends on the delay and pulse area of the third pulse, aligning with theoretical model derived by assuming excitation with $\delta-$function pulses:
@@
The optimal phase cycling methods for the signals corresponding to $\phi_{F}$ and $\phi_{S}$ were $3\times 6\times 3\times 1$ and $3\times 5\times 4\times 1$, respectively. However, we used the $3\times 6\times 4\times 1$ phase cycling method, which allowed us to isolate both signals simultaneously using the same phase combination of excitation pulses. This phase-cycling scheme is the optimal choice for our purpose. As an added advantage, using the same phase-cycling scheme for both the FWM and SWM signals allows a direct comparison of their amplitudes. This approach is also relevant for experimental settings where the same phase combination can be used to simultaneously measure both signals. We model a single QD as a 2LS and used the OBEs incorporating all decay terms similar to those in the discussed study. Our simulation results, depicted in Fig. 8 for $\theta_{3}=\pi$, show similar observations to those described in the study. During the time interval between the arrival of the second and third excitation pulses, the FWM signal decays with the population decay rate (0.01 meV) while the SWM signal amplitude remains zero. Consistent with the analytical calculations, we observed that as the third pulse arrived ($\tau_{3}>0$ ps), the total FWM signal is converted to the SWM signal, and the FWM signal value dropped to zero.
@@ IV.3 Photon echo from high-intensity excitation
As a final example, we look at a study that investigated the effect of high-intensity excitation pulses on the photon-echo transients in an ensemble of (In,Ga)As/GaAs QDs Poltavtsev et al. (2016). In a noncollinear two-pulse FWM experiment, the photon-echo signal peaks after the second excitation pulse at a delay equal to the delay between the excitation pulses. In Ref. Poltavtsev et al., 2016, experiments performed with high-intensity excitation pulses show that the photon-echo transients do not necessarily follow this behavior. Solutions of the Lindblad equation assuming rectangular temporal profile of the excitation pulses were used to model the experimental findings and infer that significant modifications in the photon-echo transients occur when the spectral width of the inhomogeneous ensemble is comparable to those of the excitation pulses.
@@
The above results explicitly demonstrate that the our calculations also work for systems with inhomogeneous broadening. As in Sec. IV.2, we did not make any simplifying assumptions about the pulse shape, as was done in the original work Poltavtsev et al. (2016). Collectively, the results presented in this section highlight the effectiveness of the phase-cycling method to calculate the exact nonlinear response even for high-intensity excitation pulses. Interestingly, the calculations in in Secs. IV.2 and IV.3 were performed for excitation pulses with pulse area $\theta\geq\pi$. At such high-intensities, perturbative calculations cannot be used even if perturbative contributions up to an arbitrarily high order are included. The fact that we can reproduce experimental results in this regime further highlights the method’s usefulness. While the formalism presented here assumes collinear pulses and population-detected signals, the experimental results reproduced in this study were based on heterodyne-detected radiated-polarization signals. This underscores the broader applicability of our approach irrespective of the detection scheme of signal used in experiments.
@@ V Conclusion
In this work we have presented a comprehensive study on the application of the phase-cycling method to simulate coherent nonlinear signal for high-intensity excitation pulses. We started with a description of the phase-cycling method for a 2LS that can be applied to an arbitrary NWM signal. Following this, we used the phase-cycling method to reproduce the experimentally-measured saturation of the FWM response of QW excitons as the excitation pulse intensity is increased. Finally, we reproduced three previously-reported experimental results on measuring coherent nonlinear response of QDs using high-intensity excitation pulses, including experiments performed in nonperturbative regimes with pulse area of excitation pulses $\sim\pi$.
@@
In addition to the results discussed in this work, coherent interaction of high-intensity pulses with two- or few-level systems can be interesting to understand several other phenomena, such as many-body interactions Meinardi et al. (2003), exciton-exciton annihilation Tiwari, Peters, and Jonas (2012), overcoming electronic dephasing and vibrational relaxation Egorova et al. (2008), and manipulating nonadiabatic couplings Chen, Bennett, and Eisfeld (2022). Beyond these, V-type systems, which naturally occur in materials such as transition metal dichalcogenides (TMDCs) Wang and Thoss (2008), semiconductor nanostructures Wang et al. (2005), and atoms Lim et al. (2011), also provide opportunities to investigate even richer dynamics. These systems feature excited states coupled through a common ground state, enabling studies of quantum interference among multiple-order signals under high-intensity conditions Wu et al. (2024); Binz et al. (2020). This work can also be relevant to understand the effects of high-intensity exciation on complex phenomena such as electron-vibrational coupling in molecular systems where the Hamiltonian can be reduced to consider a few eigen states Lewis and Ogilvie (2012); Duan et al. (2017).
Finally, we address some limitations of the simulations performed in this work, which are primarilty related to the validity of the OBEs and the RWA. For instance, the OBEs are not valid in case of non-Markovian decay dynamics Berman (1986) and also do not include multiphoton excitation terms that may be relevant for high-intensity excitation Kay (1981). Similarly, the RWA will not be valid for high-intensity excitation as the Rabi frequency $\Omega_{R}$ approaches or exceeds the pulse carrier frequency $\omega_{L}$ Berman (2011). Our simulations align well with the experimental results discussed, confirming the validity of the RWA under these conditions. We also note that the phase-matching condition is a manifestation of the RWA; the wave-mixing experiments, as are commonly interpretated, would not work if the RWA is not valid. Nevertheless, it is instructive to estimate the region of validity of the simulations from the perspective of the RWA. While the Rabi frequency is typically defined for continuous-wave (CW) lasers, it is not straightforward to define for pulsed lasers Allen and Eberly (1987). However, analogous to CW excitation, we can define average Rabi frequency $\langle\Omega_{R}\rangle=\theta/\Delta t$ for pulsed excitation in terms of the pulse area $\theta$ and FWHM $\Delta t$. Using this approximation, for an optical laser pulse centered at $\omega_{L}=800$ nm with $\Delta t=100$ fs and $\theta=2\pi$, we obtain the ratio $\omega_{L}/\langle\Omega_{R}\rangle\approx 40$. Thus, the RWA and, consequently, our calculations should hold even for the above situation, which assumes a significantly higher $\langle\Omega_{R}\rangle$ compared to those used in this work.
Appendix A Numerical Simulation Framework
Our simulation framework begins with the reformulation of the OBEs in the rotating frame of reference, utilizing the RWA to retain only resonant terms, as exemplified for a two-level system by:
${\dot{\rho}^{{}^{\prime}}}_{11}$ $=-\Gamma_{1}\rho^{{}^{\prime}}_{11}-\frac{i\mu G_{tot}(t)}{2\hbar}(\rho^{{}^{\prime}}_{01}-\rho^{{}^{\prime}*}_{01})$ (13a)
${\dot{\rho}^{{}^{\prime}}}_{01}$ $=-(i\Delta_{01}+\gamma_{01})\rho^{{}^{\prime}}_{01}+\frac{i\mu G_{tot}(t)}{2\hbar}(\rho^{{}^{\prime}}_{00}-\rho^{{}^{\prime}}_{11}),$ (13b)
Here, $\rho^{{}^{\prime}}_{ij}$ = $\rho_{ij}e^{-i\omega_{L}t}$ and $\rho^{{}^{\prime}}_{ii}$ = $\rho_{ii}$ represent the redefined matrix elements in the rotating frame. The total electric field amplitude is given by $G_{tot}=\sum_{j=1}^{M}G_{j}(t-t_{j})$, and the detuning is defined as $\Delta_{01}=\omega_{01}-\omega_{L}$, where $\omega_{L}$ is the center frequency of the excitation pulse. Each complex density matrix element is separated into real and imaginary parts.
As illustrated in Fig. 1(b), the pulse sequence is defined with arrival times ($t_{j}$) for each pulse. The first pulse ($t_{1}$) can theoretically be set to zero; however, this would result in half of the pulse lying in the negative time domain. To avoid solving OBEs for negative time, is shifted to a positive time value, typically a few times the pulse duration ($\Delta t$). For instance, in the IV.3 case, with $\Delta t$ = 2.9 ps, $t_{1}$ was set at 15 ps. Arrival times for subsequent pulses ($t_{j}$) are defined recursively as $t_{j}=t_{1}+\tau_{1}+....+\tau_{j-1}$. The number of scanned or fixed delays can be adjusted based on the desired nonlinear signal and experimental dimensionality. For example, in rephasing FWM signals, $\tau_{1}$ and $\tau_{3}$ are scanned while $\tau_{2}$ remains fixed.
Numerical integration is performed using Python’s odeint solver from Scipy library, with the time step set significantly smaller than the pulse duration to accurately resolve pulse overlap effects. For a 2.9 ps pulse, a time step of 0.25 ps was used. The integration time range is defined as $t_{total}=2t_{1}+\tau_{1}+.....+\tau_{M-1}$ where the additional $t_{1}$ ensures no overlap between the last pulse and the detection time. Initial conditions are set such that the system starts in the ground state, with no population in the excited state or coherence terms.
Nested loops are implemented to scan parameters such as delays, electric field strengths, and phase cycling steps. The computations are parallelized using Python’s multiprocessing library to optimize efficiency and reduce runtime. This framework ensures accuracy in capturing system dynamics while maintaining computational feasibility.
Appendix B Optimized phase-cycling scheme